OpenStudy (anonymous):

prove that 1-cos5x= 1-(1-sin^2 5x/2)

OpenStudy (anonymous):

any one answer this please

OpenStudy (anonymous):

Note that, by the cosine subtraction formula: cos(A - B) = cos(A)cos(B) + sin(A)sin(B). So, we have: LHS = 1 - cos(5x)cos(3x) - sin(5x)sin(3x) = 1 - [cos(5x)cos(3x) + sin(5x)sin(3x)], by factoring out the negative = 1 - cos(5x - 3x), by the above = 1 - cos(2x) = 1 - [1 - 2sin^2(x)], by the double angle formula for cosine = 2sin^2(x) = RHS.